# Instance-dependent uniform tail bounds for empirical processes

### 07 Feb 2024 v4

@@ 4.1 Generalized $\gamma$ functional
Evaluating or even finding a good approximation for a $\gamma$ functional of a general set $\mathcal{A}$ can be challenging [Tal01, vHan18a], and the only solution could be “guessing” an appropriate majorizing measure or an admissible sequence of subsets [Tal01]. By pulling the supremum into the summation in the definition of $\gamma_{\alpha}$ functional, the infimum over the admissible sets would be achieved with each $\mathcal{A}_{i}$ being a covering set of $\mathcal{A}$ of cardinality $2^{2^{i}}$. This approximation describes the Dudley’s (entropy) integral inequality (see, e.g., [Ver18, Theorem 8.1.3], [Dir15, equation 2.3]), i.e.,
@@
where $N(\mathcal{A},\rho,\varepsilon)$ is the covering number of $\mathcal{A}$ with respect to $\rho$-balls of radius $\varepsilon$, and $\lesssim_{\alpha}$ is the usual inequality sign up to a (positive) constant factor depending only on $\alpha$. If accurate estimates of the covering numbers of $\mathcal{A}$ are available, approximations of $\gamma_{\alpha}$ through Dudley’s inequality are easy to compute. However, Dudley’s inequality may not deliver sufficiently sharp approximations (see, e.g., [vHan18a, Section 3.1]). The notable approach of [vHan18a, ] improves on Dudley’s inequality by replacing the entropy numbers of the entire set $\mathcal{A}$ by those of certain scale-dependent “thin” subsets of $\mathcal{A}$, imitating the multiscale form of $\gamma_{\alpha}$. These thin subsets are “smoothed projections” of $\mathcal{A}$ expressed by minimizers of interpolation of the base metric and a given nonnegative functional at different scales [vHan18a, Section 2.1]. The resulting approximation of $\gamma_{\alpha}$ is shown to be sharp in several nontrivial examples where Dudley’s inequality yields rather loose approximations [vHan18a, Section 3].

### 21 Mar 2024 v5

@@ 4.1 Generalized $\gamma$ functional
Evaluating or even finding a good approximation for a $\gamma$ functional of a general set $\mathcal{A}$ can be challenging [Tal01, vHan18], and the only solution could be “guessing” an appropriate majorizing measure or an admissible sequence of subsets [Tal01]. By pulling the supremum into the summation in the definition of $\gamma_{\alpha}$ functional, the infimum over the admissible sets would be achieved with each $\mathcal{A}_{i}$ being a covering set of $\mathcal{A}$ of cardinality $2^{2^{i}}$. This approximation describes the Dudley’s (entropy) integral inequality (see, e.g., [Ver18, Theorem 8.1.3], [Dir15, equation 2.3]), i.e.,
@@
where $N(\mathcal{A},\rho,\varepsilon)$ is the covering number of $\mathcal{A}$ with respect to $\rho$-balls of radius $\varepsilon$, and $\lesssim_{\alpha}$ is the usual inequality sign up to a (positive) constant factor depending only on $\alpha$. If accurate estimates of the covering numbers of $\mathcal{A}$ are available, approximations of $\gamma_{\alpha}$ through Dudley’s inequality are easy to compute. However, Dudley’s inequality may not deliver sufficiently sharp approximations (see, e.g., [vHan18, Section 3.1]). The notable approach of [vHan18, ] improves on Dudley’s inequality by replacing the entropy numbers of the entire set $\mathcal{A}$ by those of certain scale-dependent “thin” subsets of $\mathcal{A}$, imitating the multiscale form of $\gamma_{\alpha}$. These thin subsets are “smoothed projections” of $\mathcal{A}$ expressed by minimizers of interpolation of the base metric and a given nonnegative functional at different scales [vHan18, Section 2.1]. The resulting approximation of $\gamma_{\alpha}$ is shown to be sharp in several nontrivial examples where Dudley’s inequality yields rather loose approximations [vHan18, Section 3].